OpenStudy (anonymous):

if log3=x and log7=y, then log63 is......

jimthompson5910 (jim_thompson5910):

Hint: 63 = 9*7 = 3^2*7

OpenStudy (anonymous):

x^2y?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

that allows us to say \[\Large \log(63) = \log(3^2*7)\]

jimthompson5910 (jim_thompson5910):

what can we do next?

OpenStudy (anonymous):

well log3 ^2 is x sooooo x *x*y

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

do you know the log rules?

OpenStudy (anonymous):

kind of......

jimthompson5910 (jim_thompson5910):

review these rules http://www.purplemath.com/modules/logrules.htm

jimthompson5910 (jim_thompson5910):

hopefully they seem familiar

OpenStudy (anonymous):

2x+y?

jimthompson5910 (jim_thompson5910):

very good

OpenStudy (anonymous):

so thats the answer ?

jimthompson5910 (jim_thompson5910):

yeah

OpenStudy (anonymous):

ok thank you so much!!!!

jimthompson5910 (jim_thompson5910):

np